OpenStudy (anonymous):

Simplify to get Tx/Ty by itself... [ x-z(Tx) ]/[ x-z(Ty) ] = 1

OpenStudy (anonymous):

\[{x-z(Tx) \over x-z(Ty)} = 1\]

OpenStudy (loser66):

what is T?

OpenStudy (anonymous):

I'll rewrite the equation. Need the answer to be T/R = something\[{x-z(T) \over x-z(R)} = 1\]

OpenStudy (loser66):